# GenUP: Generative User Profilers as In-Context Learners for Next POI Recommender Systems

### v2

@@ 5.3.1. Impact of Natural Language User Profile Components
No user profile: Baseline method of LLM4POI* with no system prompts (Li et al., 2024).
@@
User Profile + Routines & Preferences: The user profile along with the predicted routines and preferences based on the Theory of Planned Behavior (Ajzen, 1991).
User Profile + Routines & Preferences + Attributes: The user profile, predicted routines and preferences, and predicted user attributes (age, gender, educational background, socioeconomic status) (Chen et al., 2024b).
@@ 5.3.2. User Cold-start Analysis
| System Prompt | Inactive | Normal | Very Active |
| 200-word User Profile | 0.2012 | 0.2737 | 0.2728 |
| + Routines & Preferences | 0.2105 | 0.2774 | 0.2680 |
| + User Attributes | 0.2198 | 0.2664 | 0.2692 |
| + BFI Traits | 0.2105 | 0.2591 | 0.2752 |
@@ 5.3.3. Efficiency Comparison of Trajectory Similarity and User Profiling
5.3.4. Open-source and Lightweight Base LLM
As shown in Appendix F, our approach demonstrates competitive results compared to other LLM-based methods, all while utilizing open-source and lightweight models. Remarkably, the GenUP-Llama-3.2-1B model exhibited impressive predictive capabilities despite its smaller size, making it a promising candidate for offline or on-device POI recommendations, such as on mobile devices. Importantly, the reported results of LLM-ZS, LLM-Mob, and AgentMove in Table 4 used the closed-source GPT-4o Mini as their base LLM. Originally, LLM-Mob utilized GPT-3.5 Turbo (Wang et al., 2023), whereas LLM-ZS requires at least a 70 billion, instruction-tuned LLM or closed-source models like GPT-3.5, GPT-4, or GPT-4o to start exhibiting significant advantages (Beneduce et al., 2024). Although we also employed GPT-4o mini as our NL user profile generator in the experiments, this step can easily be replaced with open-source alternatives, such as instruction-tuned variants of the Llama 3.2 models that offer similar generation capabilities.
@@ 6. Conclusion and Future Work
In this paper, we propose a novel approach for next point-of-interest (POI) recommendation by introducing natural language (NL) user profiles generated from location-based social network (LBSN) check-in records. Our method addresses key challenges in existing POI recommendation systems, such as cold-start issues and the lack of interpretability in user embeddings. By leveraging natural language user profiles, we aim to improve the personalization of recommendations, reduce computational complexity, and enhance transparency and user understanding of the recommendation process.
Our methodology consists of generating NL user profiles by analyzing users’ LBSN check-ins. We incorporate behavioral theories and user attributes to create a more grounded and comprehensive profile, capturing the user’s routines, preferences, and behavioral patterns. This approach eliminates the need for long historical trajectory prompts, allowing for more efficient context lengths in LLM-based recommendations. To validate our method, we conducted experiments on multiple LBSN datasets, comparing the performance of GenUP models with baseline methods, including models using long historical trajectories and agentic LLM frameworks. The results demonstrate that our approach outperforms traditional user embedding methods and agentic LLM frameworks, particularly in terms of personalization and efficiency. The analysis of these results highlights the benefits of using interpretable NL profiles, which provide a clearer understanding of user preferences and allow for easier adaptation and personalization in real-world applications. This makes our approach a promising solution for scalable, efficient, and transparent POI recommendation systems.
@@ Acknowledgements.
Appendix A Sample User Profile Generation Results
We provide a sample generation result of a user profile from the NYC dataset in Figure 2. We used GPT-4o Mini to generate profiles given the check-in records of each user.
Appendix B User Profile Validation
We followed (Ramos et al., 2024) and their qualitative case study on NL profiles, adopting their metrics to assess our generated user profiles. LLMs were prompted using the validation prompt shown in Table 8, with evaluation results presented in Table 9.
| prompt | Given the following natural language (NL) user profile [user profile], you are to assess the user profile based on four criteria: |
| - Fluency: Is the NL profile both syntactically and semantically correct? | |
| - Informativeness: Does the NL profile provide important information for a user profile? | |
| - Conciseness: Is the NL profile written in a concise manner? | |
| - Relevance: Given the list of reviews, is the NL profile relevant to the user? | |
| Metric | NYC | CA | TKY |
| Judge: GPT 4-o Mini | |||
| Fluency | 100.00 | 100.00 | 100.00 |
| Informativeness | 100.00 | 100.00 | 100.00 |
| Conciseness | 43.17 | 37.26 | 40.60 |
| Relevance | 100.00 | 100.00 | 100.00 |
| Judge: Gemini 1.5 Flash | |||
| Fluency | 100.00 | 100.00 | 100.00 |
| Informativeness | 100.00 | 100.00 | 100.00 |
| Conciseness | 65.81 | 57.68 | 63.48 |
| Relevance | 100.00 | 100.00 | 100.00 |
Appendix C User Profile Distribution
We present the distribution of the generated NL user profiles across three datasets: NYC, CA, and TKY. Table 10 displays the distribution of predicted personality traits, while Table 11 shows the distribution of user attributes, highlighting the diversity within each dataset.
| NYC | CA | TKY | |
| Extroverted | 736 | 3526 | 1476 |
| Introverted | 311 | 430 | 805 |
| Agreeable | 1047 | 3956 | 2281 |
| Antagonistic | 0 | 0 | 0 |
| Conscientious | 1047 | 3956 | 2281 |
| Unconscientious | 0 | 0 | 0 |
| Neurotic | 41 | 268 | 99 |
| Emotionally Stable | 1006 | 3688 | 2182 |
| Open to Experience | 1043 | 3947 | 2279 |
| Closed to Experience | 4 | 9 | 2 |
| NYC | CA | TKY | ||
| Age | Child | 1 | 14 | 2 |
| Adolescent | 72 | 330 | 274 | |
| Adult | 974 | 3612 | 2005 | |
| Older Adult | 0 | 0 | 0 | |
| Gender | Male | 774 | 2911 | 1881 |
| Female | 273 | 1045 | 400 | |
| Education | Some Schooling | 43 | 190 | 192 |
| High School | 55 | 176 | 165 | |
| College & Beyond | 949 | 3590 | 1924 | |
| Socioeconomic | Lower | 3 | 7 | 3 |
| Middle | 1028 | 3651 | 2240 | |
| Upper | 16 | 298 | 38 |
Appendix D Dataset Statistics
We present the detailed statistics of our pre-processed version of the Global-scale Check-in Dataset in Table 12. AgentMove used an unspecified minimum distance to filter POI locations, whereas we used real geometric boundaries of cities to filter them.
| City | Users | Trajectories | POIs | Check-ins |
| AgentMove (Feng et al., 2024) | ||||
| Moscow | 10,501 | 100,854 | 93,599 | 950,898 |
| São Paulo | 11,856 | 77,120 | 78,904 | 809,198 |
| Ours | ||||
| Moscow | 6,864 | 66,216 | 89,247 | 500,213 |
| São Paulo | 3,718 | 40,045 | 64,690 | 300,044 |
Appendix E Hyperparameters
We show the different training hyperparameters of our LLMs across different datasets in Table 13. We also show the LoRA parameters we used for parameter-efficient fine-tuning in Table 14.
| Model | Learning | Batch Size | |
| Rate | NYC | Others | |
| GenUP-Llama-2-7B | 2e-5 | 1 | 2 |
| GenUP-Llama-3.1-8B | 2e-4 | 1 | 2 |
| GenUP-Llama-3.2-1B | 2e-4 | 1 | 2 |
| parameter | value |
| rank | 8 |
| alpha | 16 |
| dropout | 0.0 |
| bias | none |
| modules | query, key, value, output |
Appendix F Computational Efficiency Comparison
In Table 15, we compare the base LLMs used in our approach with those employed by existing methods. Our method leverages open-source, lightweight models, while previous works such as LLM-ZS (Beneduce et al., 2024), LLM-Mob (Wang et al., 2023), and AgentMove (Feng et al., 2024) rely on larger and/or or closed-source LLMs.
| Method | Base Model | #params | NYC | TKY | Open |
| LLM-ZS♯ | Llama 2 Chat | 7B | 0.086 | 0.047 | ✓ |
| LLM-ZS♯ | Llama 2 | 7B | 0.009 | 0.005 | ✓ |
| LLM-ZS♯ | Mistral | 7B | 0.087 | 0.069 | ✓ |
| LLM-ZS♯ | Llama 3 | 8B | 0.011 | 0.009 | ✓ |
| LLM-ZS♯ | Llama 3 Instruct | 8B | 0.101 | 0.081 | ✓ |
| LLM-ZS♯ | Llama 3.1 | 8B | 0.019 | 0.01 | ✓ |
| LLM-ZS♯ | Llama 2 Chat | 13B | 0.099 | 0.078 | ✓ |
| LLM-ZS♯ | Llama 2 | 13B | 0.011 | 0.009 | ✓ |
| LLM-ZS♯ | Llama 2 Chat | 70B | 0.179 | 0.199 | ✓ |
| LLM-ZS♯ | Llama 2 | 70B | 0.062 | 0.042 | ✓ |
| LLM-ZS♯ | Llama 3 | 70B | 0.084 | 0.077 | ✓ |
| LLM-ZS♯ | Llama 3 Instruct | 70B | 0.196 | 0.199 | ✓ |
| LLM-Mob♭ | GPT-3.5 Turbo | N/A | 0.274 | N/A | |
| LLM-ZS♯ | GPT-3.5 Turbo | N/A | 0.192 | 0.199 | |
| LLM-ZS♯ | GPT-4 | N/A | 0.195 | 0.205 | |
| LLM-ZS♯ | GPT-4o | N/A | 0.199 | 0.209 | |
| LLM-ZS† | GPT-4o Mini | N/A | N/A | 0.175 | |
| LLM-Mob† | GPT-4o Mini | N/A | N/A | 0.175 | |
| AgentMove† | GPT-4o Mini | N/A | N/A | 0.185 | |
| GenUP | Llama 3.2 | 1B | 0.248 | 0.185 | ✓ |
| GenUP | Llama 2 | 7B | 0.257 | 0.169 | ✓ |
| GenUP | Llama 3.1 | 8B | 0.258 | 0.212 | ✓ |
[…]

### v3

@@ 5.3.1. Impact of Natural Language User Profile Components
No user profile: Baseline method of LLM4POI* with no system prompts.
@@
User Profile + Routines & Preferences: The user profile along with the predicted routines and preferences based on the Theory of Planned Behavior.
User Profile + Routines & Preferences + Attributes: The user profile, predicted routines and preferences, and predicted user attributes (age, gender, educational background, socioeconomic status).
User Profile + Routines & Preferences + Attributes + BFI Traits: The full set, including the user profile, predicted routines and preferences, predicted user attributes, and predicted Big Five Inventory traits.
@@ 5.3.2. User Cold-start Analysis
| System Prompt | Inactive | Normal | Very Active |
| 200-word User Profile | 0.2012 | 0.2737 | 0.2728 |
| + Routines & Preferences | 0.2105 | 0.2774 | 0.2680 |
| + User Attributes | 0.2198 | 0.2664 | 0.2692 |
| + BFI Traits | 0.2105 | 0.2591 | 0.2752 |
@@ 5.3.3. Efficiency Comparison of Trajectory Similarity and User Profiling
| Target Category | Original Profile | Added New Preference |
| Coffee Shop | 61 | 68 (+7) |
| Bar | 94 | 105 (+11) |
| Gym / Fitness Center | 102 | 123 (+21) |
| Subway | 53 | 65 (+12) |
5.3.4. Preference Shift Adaptability
To evaluate the responsiveness of GenUP to short-term user preference changes, we conducted an experiment on the NYC test set using the GenUP-Llama-3.2-1B model. Following the approach of prior work (Ramos et al., 2024), we simulate short-term shifts in user preference by appending a new preference statement to the initial NL user profile of all users. We added the prompt "Today, this user really wants to visit a {POI category} place." to the NL user profile and repeated the evaluation process without retraining the model. We selected four commonly visited POI categories for evaluation: Coffee Shop, Bar, Gym/Fitness Center, and Subway Station. For each category, we measured the number of predicted visits to that POI category before and after injecting the preference prompt into the user’s profile. The evaluation focuses on the relative change in predicted category frequencies, with the expectation that predictions will shift toward the injected preference.
Naturally, the degree of change depends on multiple contextual factors, including the user’s historical affinity for the target category and the temporal feasibility of such visits (e.g., bar visits in the morning are unlikely). Despite these constraints, this setting serves as a meaningful diagnostic for adaptability. Results in Table 12 indicate that GenUP successfully increases the predicted frequency of the target category across all four target categories, demonstrating its ability to promptly adapt its predictions in response to NL user profile edits. Crucially, this flexibility is achieved without any model fine-tuning, highlighting the efficiency of the NL user profile prompting framework. These findings underscore the potential of GenUP for real-time preference adaptation, enabling dynamic, user-controllable personalization in next POI recommendation systems.
Example from the Moscow dataset where both GenUP and LLM4POI* incorrectly predicted a novel POI instead of the ground-truth revisit to a previously visited location within the same trajectory.
Example from the São Paulo dataset illustrating a failure case where both GenUP and LLM4POI* were overly conservative, predicting historically visited POIs instead of a plausible but novel ground-truth destination.
5.3.5. Qualitative Analysis
To further investigate how GenUP compares to baseline models lacking user profiles, we conducted a qualitative analysis on the Moscow and São Paulo datasets. Specifically, we examined two representative cases: (1) instances where both GenUP and the LLM4POI* baseline incorrectly predicted the next POI, and (2) instances where all GenUP variants correctly predicted the next POI while LLM4POI* failed.
In the first scenario, we observed two types of systematic errors. In some cases, both GenUP and LLM4POI* were overly explorative, predicting a novel POI instead of returning to a previously visited location within the same trajectory. Figure 3 illustrates such a case from the Moscow dataset. Conversely, another failure mode emerged when the models were too conservative, overly reliant on historical preferences, and failed to explore plausible but novel POIs. This behavior is depicted in a São Paulo case shown in Figure 4. These observations highlight the inherent tension and tradeoff between exploiting routine patterns and exploring new but contextually relevant next POI options.
Examples where all GenUP models correctly predicted the next POI, while LLM4POI did not. In both (a) Moscow and (b) São Paulo, the user returns to a previously visited POI, correctly captured by GenUP.
In the second scenario, where all GenUP models correctly predicted the next POI and LLM4POI* failed, both models generated predictions that correspond to previously visited POIs within the current trajectory. However, only GenUP correctly identified the target POI that the user revisited. As shown in Figure 5(b), which presents representative cases from the Moscow and São Paulo datasets, GenUP disambiguates among multiple historically plausible POIs by leveraging behavioral patterns contained in the natural language user profile. Whereas LLM4POI* operates solely on short-term trajectory context, GenUP conditions its predictions on a richer representation of long-term mobility behaviors. These examples highlight GenUP’s advantage in routine-driven prediction scenarios, where personalized profiles inform the model with patterns and routines not found in the current trajectory alone.
@@ 6. Conclusion and Future Work
In this paper, we propose a novel framework for next POI recommendation by introducing generative natural language user profiles derived from LBSN check-in sequences. To the best of our knowledge, this is the first approach that generates NL user profiles directly from spatio-temporal, check-in data and integrates them into a next POI prediction framework. Our method addresses key challenges in POI recommendation, including cold-start scenarios and the lack of interpretability in embedding-based systems. By incorporating behavioral theories and user attributes, our profiles capture routines, preferences, and long-term mobility patterns to inform personalized POI recommendations. Our results demonstrate that our approach outperforms existing LLM-based baselines in terms of personalization, efficiency, and adaptability. Specifically, we observe consistent improvements in Acc@1 across multiple datasets, underscoring GenUP’s ability to better capture user-specific preferences. Our method achieves greater computational efficiency by eliminating the need for long input contexts and costly trajectory similarity computations required by prior work, such as LLM4POI. Moreover, we show that GenUP adapts to user preference shifts, where profiles can be updated with simple natural language edits without model re-training. Qualitative analyses further reveal that GenUP captures long-term user routines more effectively, enabling more accurate and context-aware predictions. Together, these findings highlight the promise of interpretable, editable NL user profiles to build more transparent and efficient LLM-based POI recommendation systems. In the future, we plan to introduce intent modeling via reasoning, which would provide greater transparency and further improve the accuracy of recommendations by more accurately capturing users’ underlying motivations and decision-making processes.